# A preconditioned difference of convex functions algorithm with extrapolation and line search

### v1

@@ 4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimal condition. Moreover, the iteration sequence $\{x^{n}\}_{n}$ is bounded.
@@
As for the framework of Algorithm 2, we can naturally obtain global convergence. Noting that the difference between Algorithm 1 and Algorithm 2 is the difference of (4.3) and (4.19). In the above proof, we have only used (4.3) after the (4.21); therefore, in the framework of Algorithm 2, the subsequent proofs are not quite different from the proofs for Algorithm 1.
@@ 5 Numerical experiments
  • SCAD-Regularized Least Squares: Implemented on an AMD Ryzen 7 8845H system (3.80 GHz CPU with Radeon 780M iGPU).
  • Nonlocal Ginzburg-Landau Segmentation: Executed on NVIDIA RTX 3080 Ti GPU system (16GB VRAM).
Detailed implementation protocols and comparative results will be systematically presented in subsequent subsections. For all tables, whence the iteration number exceeds the maximum iterations, we use notation — instead of recording the CPU time.
@@ 5.1 Least squares problems with modified SCAD regularizer
[…]
Choosing $M=\lambda_{A^{T}A}I-A^{T}A$, we can see that the two algorithms have the same iteration.
@@
Precision Iter CPU time (s)
@@
Precision Iter CPU time (s)
@@
$10^{-4}$ 202 40967 1216 193 85 0.8 130.6 4.0 1.8 0.8
$10^{-5}$ 1002 131326 20644 4375 676 3.4 422.1 67.8 47.8 5.4
$10^{-6}$ 8202 Max 168123 27154 2905 27.6 528.4 257.1 21.1
$10^{-7}$ 39002 Max Max 192202 10118 127.6 1793.4 76.3
$10^{-8}$ 135802 Max Max Max 11745 428.8 93.1
$10^{-9}$ 451002 Max Max Max 14298 1458.7 113.2
(a) $l_{1}$-convergence-rate
(b) Huber-convergence-rate
@@ 5.2 Graphic Ginzburg-Landau model for image segmentation
where $x$ denotes segmentation labels, $w_{ij}=K(i,j)\cdot N(i,j)$ encodes nonlocal interactions via Gaussian similarity $K(i,j)=\exp\left(-\|P_{i}-P_{j}\|_{2}^{2}/\kappa^{2}\right)$ and binary adjacency $N(i,j)$, $\mathbb{W}(x)=\frac{1}{4}\sum_{i=1}^{N}(x(i)^{2}-1)^{2}$ enforces binary constraints, the diagonal matrix $\Lambda$ and the vector $y$ incorporate prior labels [28]. We choose a $25\times 25$ square box surrounding each pixel for the nonlocal interactions. Parameters $\tau=\gamma=10$ balance regularization strength.
The above energy function $E(x)$ is divided into two components
@@
The first criterion “DICE Bound” in Table 4 relies on the segmentation coefficient DICE, a metric used to evaluate the segmentation outcome. It is calculated using the formula $\text{DICE}=\frac{2|X\cap Y|}{|X|+|Y|}$, where $|X|$ and $|Y|$ denote the pixel counts in the segmentation and ground truth, respectively. Moreover, the criterion “DICE Bound” refers to the DICE coefficient of the segmentation achieving a value of $0.985$, indicating a high degree of similarity.
@@ 6 Conclusion
Acknowledgements Ran Zhang and Hongpeng Sun acknowledge the support of the National Natural Science Foundation of China under grant No.  12271521, National Key R&D Program of China (2022ZD0116800), and Beijing Natural Science Foundation No. Z210001. The second author acknowledges Prof. Ting Kei Pong of The Hong Kong Polytechnic University for suggesting the techniques developed in [22] during a private communication.

### v2

@@ 4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimality condition. Moreover, the iteration sequence $\{x^{n}\}_{n}$ is bounded.
@@
As for the framework of Algorithm 2, we can naturally obtain global convergence. Noting that the difference between Algorithm 1 and Algorithm 2 is the difference of (4.4) and (4.20). Therefore, in the framework of Algorithm 2, the subsequent proofs are not quite different from the proofs for Algorithm 1.
@@ 5 Numerical experiments
  • SCAD-Regularized Least Squares: Implemented an NVIDIA RTX 4070 GPU system (24GB VRAM).
  • Nonlocal Ginzburg-Landau Segmentation: Executed on an NVIDIA RTX 3080 Ti GPU system (16GB VRAM).
Detailed implementation protocols and comparative results will be systematically presented in subsequent subsections. For all tables, when the iteration number exceeds the maximum iterations, we use the notation “—” instead of recording the CPU time.
@@ 5.1 Least squares problems with modified SCAD regularizer
$f(x)=\frac{1}{2}\|Ax-b\|^{2},\ g_{1}(x)=\mu\|x\|_{1}\ \text{ and }\ g_{2}(x)=\mu\sum_{i=1}^{k}\int_{0}^{|x_{i}|}\frac{[\min\{\theta\mu,u\}-\mu]_{+}}{(\theta-1)\mu}du.$
Choosing $M=\lambda_{A^{T}A}I-A^{T}A$, we can see that the two algorithms have the same iterations.
@@
Remark 6.
In addition, we introduce an additional criterion to demonstrate the accuracy of the first-order optimality condition of the problem (5.1) and (5.2). For (5.1), the following criterion is used to measure the residual of the first-order optimality condition for the nonsmooth problem (5.1)
$res\left(x^{n}\right)=\left\|G_{\operatorname{proj}}\left(x^{n}\right)\right\|,\quad G_{\operatorname{proj}}\left(x^{n}\right)=\left(x^{n}-\operatorname{prox}_{I_{2}}\left(x^{n}-\nabla I_{1}\left(x^{n}\right)\right)\right),$
where $I_{1}\left(x\right)=\frac{1}{2}\left\|Ax-b\right\|^{2}+p_{2}\left(x\right)$, $I_{2}\left(x\right)=\mu\left\|x\right\|_{1}$, and $\operatorname{prox}_{I_{2}}(z):=\operatorname*{arg\,min}_{x}\{I_{2}(x)+\frac{1}{2}\|x-z\|^{2}$}. For (5.2), we choose $res\left(x^{n}\right):=\|\nabla E(x^{n})\|$ with $E(x)$ as in (5.2).
@@
Precision Iter GPU time (s)
@@
Precision Iter GPU time (s)
@@
$10^{-4}$ 202 42804 1216 187 100 0.5 35.0 1.4 0.6 0.3
$10^{-5}$ 1002 161030 20132 4243 506 1.5 131.9 16.8 16.1 1.5
$10^{-6}$ 8002 Max 163751 28232 2805 8.2 144.2 98.2 7.8
$10^{-7}$ 38402 Max Max 195804 9672 37.6 593.4 26.7
$10^{-8}$ 132802 Max Max Max 13545 127.8 35.4
$10^{-9}$ 437402 Max Max Max 17067 473.2 50.2
(a) $l_{1}$-Convergence-Rate (GPU)
(b) Huber-Convergence-Rate (GPU)
@@ 5.2 Graphic Ginzburg-Landau model for image segmentation
where $x$ denotes segmentation labels, $w_{ij}=K(i,j)\cdot N(i,j)$ encodes nonlocal interactions via Gaussian similarity $K(i,j)=\exp\left(-\|P_{i}-P_{j}\|_{2}^{2}/\kappa^{2}\right)$ and binary adjacency $N(i,j)$, $\mathbb{W}(x)=\frac{1}{4}\sum_{i=1}^{N}(x(i)^{2}-1)^{2}$ enforces binary constraints, the diagonal matrix $\Lambda$ and the vector $y$ incorporate prior labels [28]. We choose a $25\times 25$ square box surrounding each pixel for nonlocal interactions. Parameters $\tau=\gamma=10$ balance regularization strength.
The energy function mentioned above $E(x)$ is divided into two components
@@
The first criterion “DICE Bound” in Table 4, relies on the segmentation coefficient DICE, a metric used to evaluate the segmentation outcome. It is calculated using the formula $\text{DICE}=\frac{2|X\cap Y|}{|X|+|Y|}$, where $|X|$ and $|Y|$ denote the pixel counts in the segmentation and ground truth, respectively. Moreover, the criterion “DICE Bound” refers to the DICE coefficient of the segmentation, achieving a value of $0.985$, indicating a high degree of similarity.
@@ 6 Conclusion
Acknowledgements Ran Zhang and Hongpeng Sun acknowledge the support of National Key R&D Program of China (2022ZD0116800), the National Natural Science Foundation of China under grant No.  12271521, and Beijing Natural Science Foundation No. Z210001. The second author acknowledges Prof. Ting Kei Pong of The Hong Kong Polytechnic University for suggesting the techniques developed in [22] during a private communication.